# Equation-of-state-informed pulse profile modeling

### v3

@@ Abstract
NICER has enabled mass–radius inferences for pulsars using pulse profile modeling (PPM), providing constraints on the equation of state (EOS) of cold, dense matter. To date, PPM and EOS inference have been carried out as two separate steps, with the former using EOS-agnostic priors. This approach has several drawbacks. Ideally, one would perform a fully hierarchical Bayesian inference where the pulse profile and EOS model parameters are jointly fit, but implementing such a framework is complex and computationally demanding. Here, we present an intermediate solution introducing an EOS-informed prior on mass-radius into the existing PPM pipeline using normalizing flows. By focusing on the parameter space consistent with certain EOSs, this approach both tightens constraints on neutron star parameters while reducing computational costs and requiring minimal additional implementation effort. We test this approach on two pulsars, PSR J0740+6620 and PSR J0437-4715, and with two EOS model families: a model based on the speed of sound inside the neutron star interior (CS) and a piecewise-polytropic (PP) model. Both EOS models implement constraints from chiral effective field theory calculations of dense matter. For both pulsar datasets, the inferred radius credible intervals are narrower than in the EOS-agnostic case, with CS favoring smaller radii and PP favoring larger radii. For PSR J0437-4715, the EOS-informed priors reveal a new, more extreme geometric mode that is statistically favored but physically questionable. Including the PPM posteriors in the subsequent EOS inference further tightens the mass-radius posteriors through the chiral effective field theory constraints. However, there is also a sensitivity to the high-density extensions, where the PP (CS) model produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
@@ I Introduction
To inform the PPM pipeline of the underlying cold, dense matter EOS prior, we use the two EOS model families considered in Rutherford et al. (2024). At nuclear densities beyond the Baym-Pethick-Sutherland (BPS) crust (Baym et al., 1971), these are based on next-to-next-to-next-to-leading-order (N3LO) chiral effective theory calculations of matter in beta equilibrium from Keller et al. (2023) up to $1.5n_{0}$, where $n_{0}=0.16\,\mathrm{fm}^{-3}$ is the nuclear saturation density. The EOS prior thus builds in knowledge of the crust and of the EOS at nuclear densities based on our modern understanding of nuclear forces. At higher densities, we use two possible model extensions, based on the speed of sound inside the neutron star interior (CS) Greif et al. (2019) or based on a piecewise-polytropic (PP) model Hebeler et al. (2013). In this work, we explore the impact of both prior EOS ensembles on the PPM-derived MMRR and geometry posteriors of PSR J0740+6620 and PSR J0437-4715, and the effects of the derived MMRR posteriors on the subsequent EOS inference analysis.
The remainder of this paper is organized as follows. In Sec. II.1, we present the ideal case of a full hierarchical Bayesian model for the simultaneous inference of neutron star and EOS properties. Section II.2 introduces our intermediate approach that incorporates EOS prior information to constrain neutron star properties, followed by the EOS inference framework using EOS-prior-informed PPM MRM–R posteriors. Sections II.3 and II.4 describe the X-PSI PPM analysis pipeline and the use of normalizing flows to model the MRM–R prior derived from the two EOS families. In Sec. III, we summarize the datasets for PSR J0740+6620 and PSR J0437–4715, and outline the prior intervals for the PP and CS EOS model parameters used in this analysis. Section IV presents the EOS-prior-informed PPM posteriors and the resulting constraints on the cold, dense matter EOS. Finally, in Sec. V, we discuss our conclusions and directions for future work.
@@ II.1 Full hierarchical bayesian model
again assuming independence across neutron stars. We note that, strictly speaking, $\left\{\theta_{k}\right\}^{K}_{k=1}$, $\left\{\varepsilon_{c,k}\right\}^{K}_{k=1}$, and $\varphi$ depend on certain fixed hyperparameters. [^4] Hyperparameters are parameters that determine the shape of the prior distribution for a given model parameter. For example, if the prior for an inclination angle follows a normal distribution, the hyperparameters would be its mean and standard deviation. However, since these hyperparameters are fixed and therefore not random variables, we have chosen not to write them down for clarity. Similarly, all distributions in the joint PPM-EOS inference should be read as conditional on the choices of PPM model (see, e.g., Riley et al., 2019; Vinciguerra et al., 2023) and on the choice of EOS model (see Hebeler et al., 2013; Greif et al., 2019, for the PP and CS models, respectively). To avoid cluttering the notation, we have omitted including these in the conditional densities as well.
@@ II.2 Intermediate step: PPM including EOS prior information and EOS inferencing
For the EOS inference, we implement the Raaijmakers et al. (2021) Bayesian framework using NEoST (v2.1.0)Raaijmakers et al. (2025b). [^5] The prerelease form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023). The released version of NEoST has been used in Rutherford et al. (2024, 2025). NEoST uses MultiNest to sample from the chosen EOS parameter space and combines constraints from both the MRM–R relations derived from PPM and the $M–\Lambda$ relations from gravitational-wave observations. In this work, however, we focus solely on incorporating EOS prior information into the PPM-derived MRM–R posteriors, and therefore omit the gravitational-wave constraints to highlight the differences between the original and new frameworks. Since the EOS model parameters considered in this work—namely PP and CS—are uniformly sampled, the EOS inference framework is identical to that used in our previous work Raaijmakers et al. (2021); Rutherford et al. (2024).
@@ II.4 Normalizing flow as EOS prior
After performing the Bayesian hyperparameter tuning, we selected the best-performing configuration without further exploring alternative setups. Because there were many configurations with similar validation losses, we chose one with a relatively short training time. This configuration—using a normalizing flow with ten transformations for both $p(R|M)$ and p(M)p(M)—performed well, achieving a test loss of 2.52 for CS and 2.14 for PP (see Fig. 3 for the training and validation loss curves) in approximately 27 minutes for both cases. Other configurations with comparable loss values required substantially longer training times. The resulting network contains two hidden layers with 32 neurons each, and ReLU [^10] ReLU (rectified linear unit) is an activation function defined as $f(x)=\text{max}(0,x)$. Activation functions introduce nonlinearity into the network, allowing the model to capture complex patterns and relationships in the data. Nair and Hinton (2010) activations following each layer. The normalizing flow is trained on 70,000 MRM–R samples for 4000 epochs, using the Adam optimizer Kingma and Ba (2017) with a learning rate of $3\times 10^{-3}$. The model is optimized by minimizing the negative log-likelihood of the MMRR samples.
Figure 3: Training (orange) and validation (blue) loss curves for the PP EOS model (top) and CS EOS model (bottom).
To evaluate, we compare the probability distribution learned by the normalizing flow to our test set using the Kullback-Leibler (KL) divergence, which provides a single number quantifying how closely the entire learned distribution approximates the target one. This results in KL divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and PP EOS prior distribution, respectively. A KL divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set. See Fig. 2 for a visual comparison between the prior model and the resultant normalizing flow.
When sampling radii conditioned on a given mass, as in the cases of PSR J0740+6620 and PSR J0437-4715, samples with radii outside the interval 5.5 to 14.0 km for the PP EOS model and 6.5 to 14.5 km for the CS EOS model are excluded from the analysis. These bounds are chosen to avoid artifacts arising in regions where the normalizing flow assigns low but nonzero probability and to avoid drawing unphysical MMRR values. The bounds are conservative and lie well outside the support of the EOS prior. For a sample size of $1\times 10^{6}$, the probability of drawing samples outside these bounds is close to zero, so truncation is not expected to significantly affect the normalization of the distribution.
@@ III Data
In this section, we outline the x-ray event data used in this study. In Sec. III.1, we describe the most important aspects of the dataset from the high-mass pulsar PSR J0740+6620, as detailed in Salmi et al. (2024a), and Hoogkamer et al. (2025a). Section III.2 gives a brief description of the observations from the nearest and brightest rotation-powered millisecond pulsar PSR J0437-4715, as detailed in Choudhury et al. (2024). In Sec. III.3, we summarize the prior distributions of the PP and CS EOS models.
@@ III.2 PSR J0437–4715
III.3 PP and CS EOS prior distributions
With the EOS-informed PPM framework in place (Sec. II.2), the next step is to compute the MMRR priors for the PP and CS EOS models. This involves constructing the EOS models across the relevant density range and defining the prior ranges that specify their distributions. Following Rutherford et al. (2024), we briefly summarize the EOS construction adopted in this work and the corresponding priors.
@@ III.3 PP and CS EOS prior distributions
For the PP model, the priors are defined by varying three polytropic indices, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, as well as two transition densities, $n_{12}$ and $n_{23}$, which connect the first to the second polytrope and the second to the third, respectively. The priors on these parameters are given by: $\Gamma_{1}\in[0,8]$, $\Gamma_{2}\in[0,8]$ $\Gamma_{3}\in[0.5,8]$, and $2n_{0}\leq n_{11}\leq n_{23}\leq 8.3n_{0}$.
For the CS model, the prior distribution is defined by parametrizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$. The resulting prior distributions for both the PP and CS models span a large range of possible EOSs. For additional details on the PP and CS EOS models, see Rutherford et al. (2024); Hebeler et al. (2013); Greif et al. (2019)
Finally, all EOS priors and posteriors are required to obey causality and the constraint that neutron stars have masses $\geq 1 MM_{\odot}$. This lower mass limit is imposed to be consistent with the lowest confirmed neutron star mass observation Martinez et al. (2015) and is also motivated by theoretical models of early neutron star evolution Strobel et al. (1999) as well as the minimum remnant masses predicted by simulations of core-collapse supernovae Janka et al. (2008); Fischer et al. (2010); Radice et al. (2017); Suwa et al. (2018); Müller et al. (2025). For completeness, the resulting MMRR prior distributions of both of the PP and CS models are shown in Fig. 2.
@@ IV.1 PPM including EOS prior information
To assess the influence of an EOS-informed MRM–R prior on the inferred properties of PSR J0740+6620 and PSR J0437-4715, we evaluate two representative EOS models as priors: CS and PP. These are compared against the baseline EOS-agnostic approach.
@@ IV.1.1 PSR J0740+6620
For PSR J0740+6620, we take the headline results of Salmi et al. (2024a), using the EOS-agnostic approach, as our baseline. The analysis of Salmi et al. (2024a), using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01, reported an equatorial radius of $12.49^{+1.28}_{-0.88}$ km and a gravitational mass of $2.073^{+0.069}_{-0.069}$ $M_{\odot}$. Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and PP models as prior information on MRM–R. Table 2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
Figure 4: Radius, compactness, and mass posterior distributions using the PSR J0740+6620 joint NICER and XMM-Newton dataset conditional on the ST-U model. Three posterior distributions are shown: the results from Salmi et al. (2024a) without an EOS-informed prior (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the PP-informed MMRR prior (blue). All inference results are obtained using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01. The marginal prior probability distribution functions for each parameter are displayed as dashed-dotted lines. The shaded regions in the diagonal panels contain the 68.3% credible interval for each parameter symmetric around the median. The contours in the off-diagonal panels contain the 68.3%, 95.4%, and 99.7% credible regions.
The radius credible intervals for all three models overlap. However, the radius inferred using the CS-informed prior is $11.94^{+0.65}_{-0.58}$ km—with the median smaller than the baseline—while the PP-informed prior yielded a slightly larger median radius of $12.51^{+0.56}_{-0.66}$ km. The results are in line with our expectation, since the CS model favors softer EOSs and the PP model favors stiffer ones. However, once several sources are included, the CS and PP differences are smaller Rutherford et al. (2024). Notably, the mass and radius credible intervals obtained with the EOS-informed priors are narrower than those obtained using the EOS-agnostic MMRR prior.
The posteriors of the geometry parameters, as shown in Fig. 10 of the Appendix, match even better for both the EOS-agnostic approach and the EOS-informed prior runs, with overlapping credible intervals for all parameters. Furthermore, the difference between the maximum likelihood and Bayesian evidence among all three models is negligible. [^11] The difference in maximum log-likelihood is approximately $0.1$ across all three cases, while the evidence differs by approximately +0.51+0.51 for the CS-informed run and +0.24+0.24 for the PP-informed run relative to the EOS-agnostic approach.
In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 21k and 22k core hours for the runs using the CS-informed and PP-informed priors, respectively.

### v5

@@ Abstract
NICER has enabled mass–radius inferences for pulsars using pulse profile modeling (PPM), providing constraints on the equation of state (EOS) of cold, dense matter. To date, PPM and EOS inference have been carried out as two separate steps, with the former using EOS-agnostic priors. This approach has several drawbacks. Ideally, one would perform a fully hierarchical Bayesian inference where the pulse profile and EOS model parameters are jointly fit, but implementing such a framework is complex and computationally demanding. Here, we present an intermediate solution introducing an EOS-informed prior on mass-radius into the existing PPM pipeline using normalizing flows. By focusing on the parameter space consistent with certain EOSs, this approach both tightens constraints on neutron star parameters while reducing computational costs and requiring minimal additional implementation effort. We test this approach on two pulsars, PSR J0740+6620 and PSR J0437-4715, and with two EOS model families: a model based on the speed of sound inside the neutron star interior (CS) and a piecewise-polytropic ($P$) model. Both EOS models implement constraints from chiral effective field theory calculations of dense matter. For both pulsar datasets, the inferred radius credible intervals are narrower than in the EOS-agnostic case, with CS favoring smaller radii and $P$ favoring larger radii. For PSR J0437-4715, the EOS-informed priors reveal a new, more extreme geometric mode that is statistically favored but physically questionable. Including the PPM posteriors in the subsequent EOS inference further tightens the mass-radius posteriors through the chiral effective field theory constraints. However, there is also a sensitivity to the high-density extensions, where the $P$ (CS) model produces a shift toward larger (smaller) radii and corresponding stiffening (softening) of the pressure–energy density relation.
@@ I Introduction
To inform the PPM pipeline of the underlying cold, dense matter EOS prior, we use the two EOS model families considered in Rutherford et al. (2024). At nuclear densities beyond the Baym-Pethick-Sutherland (BPS) crust (Baym et al., 1971), these are based on next-to-next-to-next-to-leading-order (N3LO) chiral effective theory calculations of matter in beta equilibrium from Keller et al. (2023) up to $1.5n_{0}$, where $n_{0}=0.16\,\mathrm{fm}^{-3}$ is the nuclear saturation density. The EOS prior thus builds in knowledge of the crust and of the EOS at nuclear densities based on our modern understanding of nuclear forces. At higher densities, we use two possible model extensions, based on the speed of sound inside the neutron star interior (CS) Greif et al. (2019) or based on a piecewise-polytropic ($P$) model Hebeler et al. (2013). In this work, we explore the impact of both prior EOS ensembles on the PPM-derived MMRR and geometry posteriors of PSR J0740+6620 and PSR J0437-4715, and the effects of the derived MMRR posteriors on the subsequent EOS inference analysis.
The remainder of this paper is organized as follows. In Sec. II.1, we present the ideal case of a full hierarchical Bayesian model for the simultaneous inference of neutron star and EOS properties. Section II.2 introduces our intermediate approach that incorporates EOS prior information to constrain neutron star properties, followed by the EOS inference framework using EOS-prior-informed PPM MRM–R posteriors. Sections II.3 and II.4 describe the X-PSI PPM analysis pipeline and the use of normalizing flows to model the MRM–R prior derived from the two EOS families. In Sec. III, we summarize the datasets for PSR J0740+6620 and PSR J0437–4715, and outline the prior intervals for the $P$ and CS EOS model parameters used in this analysis. Section IV presents the EOS-prior-informed PPM posteriors and the resulting constraints on the cold, dense matter EOS. Finally, in Sec. V, we discuss our conclusions and directions for future work.
@@ II.1 Full hierarchical bayesian model
again assuming independence across neutron stars. We note that, strictly speaking, $\left\{\theta_{k}\right\}^{K}_{k=1}$, $\left\{\varepsilon_{c,k}\right\}^{K}_{k=1}$, and $\varphi$ depend on certain fixed hyperparameters. [^4] Hyperparameters are parameters that determine the shape of the prior distribution for a given model parameter. For example, if the prior for an inclination angle follows a normal distribution, the hyperparameters would be its mean and standard deviation. However, since these hyperparameters are fixed and therefore not random variables, we have chosen not to write them down for clarity. Similarly, all distributions in the joint PPM-EOS inference should be read as conditional on the choices of PPM model (see, e.g., Riley et al., 2019; Vinciguerra et al., 2023) and on the choice of EOS model (see Hebeler et al., 2013; Greif et al., 2019, for the $P$ and CS models, respectively). To avoid cluttering the notation, we have omitted including these in the conditional densities as well.
@@ II.2 Intermediate step: PPM including EOS prior information and EOS inferencing
For the EOS inference, we implement the Raaijmakers et al. (2021) Bayesian framework using NEoST (v2.1.0)Raaijmakers et al. (2025b). [^5] The prerelease form of NEoST was used in (Greif et al., 2019; Raaijmakers et al., 2019, 2020, 2021; Rutherford et al., 2023). The released version of NEoST has been used in Rutherford et al. (2024, 2025). NEoST uses MultiNest to sample from the chosen EOS parameter space and combines constraints from both the MRM–R relations derived from PPM and the $M–\Lambda$ relations from gravitational-wave observations. In this work, however, we focus solely on incorporating EOS prior information into the PPM-derived MRM–R posteriors, and therefore omit the gravitational-wave constraints to highlight the differences between the original and new frameworks. Since the EOS model parameters considered in this work—namely $P$ and CS—are uniformly sampled, the EOS inference framework is identical to that used in our previous work Raaijmakers et al. (2021); Rutherford et al. (2024).
@@ II.4 Normalizing flow as EOS prior
After performing the Bayesian hyperparameter tuning, we selected the best-performing configuration without further exploring alternative setups. Because there were many configurations with similar validation losses, we chose one with a relatively short training time. This configuration—using a normalizing flow with ten transformations for both $p(R|M)$ and p(M)p(M)—performed well, achieving a test loss of 2.52 for CS and 2.14 for $P$ (see Fig. 3 for the training and validation loss curves) in approximately 27 minutes for both cases. Other configurations with comparable loss values required substantially longer training times. The resulting network contains two hidden layers with 32 neurons each, and ReLU [^10] ReLU (rectified linear unit) is an activation function defined as $f(x)=\text{max}(0,x)$. Activation functions introduce nonlinearity into the network, allowing the model to capture complex patterns and relationships in the data. Nair and Hinton (2010) activations following each layer. The normalizing flow is trained on 70,000 MRM–R samples for 4000 epochs, using the Adam optimizer Kingma and Ba (2017) with a learning rate of $3\times 10^{-3}$. The model is optimized by minimizing the negative log-likelihood of the MMRR samples.
Figure 3: Training (orange) and validation (blue) loss curves for the $P$ EOS model (top) and CS EOS model (bottom).
To evaluate, we compare the probability distribution learned by the normalizing flow to our test set using the Kullback-Leibler (KL) divergence, which provides a single number quantifying how closely the entire learned distribution approximates the target one. This results in KL divergences of 0.0075 and 0.0023 for the normalizing flow of the CS and $P$ EOS prior distribution, respectively. A KL divergence close to zero indicates that the distribution learned by the normalizing flow is nearly indistinguishable from the test set. See Fig. 2 for a visual comparison between the prior model and the resultant normalizing flow.
When sampling radii conditioned on a given mass, as in the cases of PSR J0740+6620 and PSR J0437-4715, samples with radii outside the interval 5.5 to 14.0 km for the $P$ EOS model and 6.5 to 14.5 km for the CS EOS model are excluded from the analysis. These bounds are chosen to avoid artifacts arising in regions where the normalizing flow assigns low but nonzero probability and to avoid drawing unphysical MMRR values. The bounds are conservative and lie well outside the support of the EOS prior. For a sample size of $1\times 10^{6}$, the probability of drawing samples outside these bounds is close to zero, so truncation is not expected to significantly affect the normalization of the distribution.
@@ III Data
In this section, we outline the x-ray event data used in this study. In Sec. III.1, we describe the most important aspects of the dataset from the high-mass pulsar PSR J0740+6620, as detailed in Salmi et al. (2024a), and Hoogkamer et al. (2025a). Section III.2 gives a brief description of the observations from the nearest and brightest rotation-powered millisecond pulsar PSR J0437-4715, as detailed in Choudhury et al. (2024). In Sec. III.3, we summarize the prior distributions of the $P$ and CS EOS models.
@@ III.2 PSR J0437–4715
III.3 $P$ and CS EOS prior distributions
With the EOS-informed PPM framework in place (Sec. II.2), the next step is to compute the MMRR priors for the $P$ and CS EOS models. This involves constructing the EOS models across the relevant density range and defining the prior ranges that specify their distributions. Following Rutherford et al. (2024), we briefly summarize the EOS construction adopted in this work and the corresponding priors.
@@ III.3 $P$ and CS EOS prior distributions
For the $P$ model, the priors are defined by varying three polytropic indices, $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, as well as two transition densities, $n_{12}$ and $n_{23}$, which connect the first to the second polytrope and the second to the third, respectively. The priors on these parameters are given by: $\Gamma_{1}\in[0,8]$, $\Gamma_{2}\in[0,8]$ $\Gamma_{3}\in[0.5,8]$, and $2n_{0}\leq n_{11}\leq n_{23}\leq 8.3n_{0}$.
For the CS model, the prior distribution is defined by parametrizing the speed of sound $c_{s}^{2}=dP/d\varepsilon$, using five parameters $a_{1}$ to $a_{5}$, which are varied over the following ranges: $a_{1}\in[0.1,1.5]$, $a_{2}\in[1.5,12]$, $a_{3}/a_{2}\in[0.05,2]$, $a_{4}\in[1.5,37]$, and $a_{5}\in[0.1,1]$. The resulting prior distributions for both the $P$ and CS models span a large range of possible EOSs. For additional details on the $P$ and CS EOS models, see Rutherford et al. (2024); Hebeler et al. (2013); Greif et al. (2019)
Finally, all EOS priors and posteriors are required to obey causality and the constraint that neutron stars have masses $\geq 1 MM_{\odot}$. This lower mass limit is imposed to be consistent with the lowest confirmed neutron star mass observation Martinez et al. (2015) and is also motivated by theoretical models of early neutron star evolution Strobel et al. (1999) as well as the minimum remnant masses predicted by simulations of core-collapse supernovae Janka et al. (2008); Fischer et al. (2010); Radice et al. (2017); Suwa et al. (2018); Müller et al. (2025). For completeness, the resulting MMRR prior distributions of both of the $P$ and CS models are shown in Fig. 2.
@@ IV.1 PPM including EOS prior information
To assess the influence of an EOS-informed MRM–R prior on the inferred properties of PSR J0740+6620 and PSR J0437-4715, we evaluate two representative EOS models as priors: CS and $P$. These are compared against the baseline EOS-agnostic approach.
@@ IV.1.1 PSR J0740+6620
For PSR J0740+6620, we take the headline results of Salmi et al. (2024a), using the EOS-agnostic approach, as our baseline. The analysis of Salmi et al. (2024a), using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01, reported an equatorial radius of $12.49^{+1.28}_{-0.88}$ km and a gravitational mass of $2.073^{+0.069}_{-0.069}$ $M_{\odot}$. Figure 4 shows the posterior distributions, as well as the ones obtained when incorporating the CS and $P$ models as prior information on MRM–R. Table 2 shows the accompanying KL divergence (in bits) between the posterior and corresponding prior for each model.
Figure 4: Radius, compactness, and mass posterior distributions using the PSR J0740+6620 joint NICER and XMM-Newton dataset conditional on the ST-U model. Three posterior distributions are shown: the results from Salmi et al. (2024a) without an EOS-informed prior (red), the results obtained with the CS-informed MMRR prior (orange), and the ones obtained with the $P$-informed MMRR prior (blue). All inference results are obtained using MultiNest with $4\times 10^{4}$ live points and a sampling efficiency of 0.01. The marginal prior probability distribution functions for each parameter are displayed as dashed-dotted lines. The shaded regions in the diagonal panels contain the 68.3% credible interval for each parameter symmetric around the median. The contours in the off-diagonal panels contain the 68.3%, 95.4%, and 99.7% credible regions.
The radius credible intervals for all three models overlap. However, the radius inferred using the CS-informed prior is $11.94^{+0.65}_{-0.58}$ km—with the median smaller than the baseline—while the $P$-informed prior yielded a slightly larger median radius of $12.51^{+0.56}_{-0.66}$ km. The results are in line with our expectation, since the CS model favors softer EOSs and the $P$ model favors stiffer ones. However, once several sources are included, the CS and $P$ differences are smaller Rutherford et al. (2024). Notably, the mass and radius credible intervals obtained with the EOS-informed priors are narrower than those obtained using the EOS-agnostic MMRR prior.
The posteriors of the geometry parameters, as shown in Fig. 10 of the Appendix, match even better for both the EOS-agnostic approach and the EOS-informed prior runs, with overlapping credible intervals for all parameters. Furthermore, the difference between the maximum likelihood and Bayesian evidence among all three models is negligible. [^11] The difference in maximum log-likelihood is approximately $0.1$ across all three cases, while the evidence differs by approximately +0.51+0.51 for the CS-informed run and +0.24+0.24 for the $P$-informed run relative to the EOS-agnostic approach.
In terms of computational cost, the EOS-agnostic inference required 84k core hours, compared to 21k and 22k core hours for the runs using the CS-informed and $P$-informed priors, respectively.